# SE-VLN: A Self-Evolving Vision-Language Navigation Framework Based on Multimodal Large Language Models

### v1

@@ Abstract
Recent advances in vision-language navigation (VLN) were mainly attributed to emerging large language models (LLMs). These methods exhibited excellent generalization capabilities in instruction understanding and task reasoning. However, they were constrained by the fixed knowledge bases and reasoning abilities of LLMs, preventing fully incorporating experiential knowledge and thus resulting in a lack of efficient evolutionary capacity. To address this, we drew inspiration from the evolution capabilities of natural agents, and proposed a self-evolving VLN framework (SE-VLN) to endow VLN agents with the ability to continuously evolve during testing. To the best of our knowledge, it was the first time that an multimodal LLM-powered self-evolving VLN framework was proposed. Specifically, SE-VLN comprised three core modules, i.e., a hierarchical memory module to transfer successful and failure cases into reusable knowledge, a retrieval-augmented thought-based reasoning module to retrieve experience and enable multi-step decision-making, and a reflection module to realize continual evolution. Comprehensive tests illustrated that the SE-VLN achieved navigation success rates of 57% and 35.2% in unseen environments, representing absolute performance improvements of 23.9% and 15.0% over current state-of-the-art methods on R2R and REVERSE datasets, respectively. Moreover, the SE-VLN showed performance improvement with increasing experience repository, elucidating its great potential as a self-evolving agent framework for VLN.
@@ Introduction
Vision-language navigation (VLN), as a key technology connecting human natural language and robot visual navigation, aims to enable embodied agents to autonomously plan paths and complete visual navigation tasks in unseen environments based on human language instructions (Anderson et al. 2018). Despite its rapid development (Chen et al. 2024b; Schumann et al. 2024; Li et al. 2024; Lin et al. 2024), especially with the help of emerging LLMs (Zhou, Hong, and Wu 2024; Chen et al. 2024a; Li, Chen, and Lin 2024; Long et al. 2024), current VLN methods lacks the advanced navigation capabilities of natural agents (e.g., bats, migratory birds, humpback whales) and shows critical shortcomings in autonomous evolution through experience. In fact, an experienced horse refines its neural circuitry through repeated journeys. Its navigation skill distills wisdom from experience and evolves autonomously through adaptation, providing a biological blueprint for existing VLN to overcome bottlenecks of ”data dependency” and ”scene generalization”.
Current VLN methods (Li, Chen, and Lin 2024; Zhou, Hong, and Wu 2024; Chen et al. 2024a) had significant limitations in terms of static experience utilization. They treated historical trajectory data as a fixed replay buffer, only using it to maintain decision consistency for the current task, and failed to extract reusable general knowledge from dynamic processes. Another drawback of existing VLN methods (Chen et al. 2024a; Zhou, Hong, and Wu 2024) lay in their limited reasoning capabilities. They relied on the static knowledge invocation of pre-trained models and failed to effectively integrate experience for multi-step decision-making. This stood in stark contrast to the biological intelligence of natural agents such as horses, which adjusted navigation paths by combining environmental memories with real-time perceptions, helping gain precise alignment between visual perceptions and language instructions when processing complex commands. Furthermore, the ”evolution” processes of existing VLN agents (Anderson et al. 2018; Lin et al. 2024, 2025) was highly dependent on manual hyperparameter tuning and model iteration, which was completely different from the mechanism which biological agents used to achieve self-evolution through natural selection and neural plasticity.
@@ A Reflection Module
As shown in Fig.4, we examined the evolution ability of the VLN framework by adjusting the number of experience entries (set to 0, 10, 30, and 50). As the number of experience increased, the overall performance of SE-VLN showed an improvement trend, i.e., OSR improved from 64.1% to 68.0%, and both SR and SPL were also improved with experience. These indicated that SE-VLN could effectively utilize past experience for continual evolution across different tasks, thereby progressively improving its navigation performance. Moreover, we observed that as the number of tasks increased, the trend of self-evolution gradually flattened out. This might have been due to the limitations of LLMs’ reasoning capabilities, and the navigation experiences would gradually tend toward homogenization.
@@
Existing methods were categorized into two types: supervised learning approaches and training-free approaches. As shown in Table 2, we evaluated the navigation performance across various scenarios using a sampled subset of 72 scenes from the R2R dataset. To ensure a fair comparison, we utilized GPT-4o as the inference model for testing both methods. Our results demonstrated that SE-VLN, significantly outperformed MapGPT, the SOTA LLM-powered VLN method. Specifically, the SE-VLN surpassed MapGPT (GPT-4o based) by 0.55 meters in NE, 12.6% in OSR, 18.1% in SR, and 24.3% in SPL. This validated the effectiveness of our approach in improving the navigation performance of LLM-powered agents without the need for training on large-scale datasets.
Furthermore, we conducted additional comparisons of the navigation performance between SE-VLN and previous models on a larger validation set that included 11 scenes with 783 trajectories. As shown in Table 3, despite the distributional differences, SE-VLN exhibited slight improvements across key metrics compared to the results from the 72-scene experiments, with further improvements of SR and SPL by 2.3% and 3.0%, respectively. This might be attributed to the fewer number of scenes in the validation set, where the experience repository might contain similar perceptual and decision experience. This result indicated that the experience repository mechanism of the SE-VLN could effectively capture the decision-making logic across tasks. When there were task patterns in new tasks that were similar to those in the experience repository, the agent could quickly transfer and reuse relevant experiences, thereby offsetting the performance loss caused by distribution differences.
@@ Conclusion
Figure 8: Performance of Different MLLMs on the R2R Sampling Subsets.
Figure 9: The prompts for the reasoning module consisted of two parts: the prefix prompt and the input prompt. These respectively provided the prerequisite information needed for the task and specified the input required at each step.

### v2

@@ Abstract
Recent advances in vision-language navigation (VLN) were mainly attributed to emerging large language models (LLMs). These methods exhibited excellent generalization capabilities in instruction understanding and task reasoning. However, they were constrained by the fixed knowledge bases and reasoning abilities of LLMs, preventing fully incorporating experiential knowledge and thus resulting in a lack of efficient evolutionary capacity. To address this, we drew inspiration from the evolution capabilities of natural agents, and proposed a self-evolving VLN framework (SE-VLN) to endow VLN agents with the ability to continuously evolve during testing. To the best of our knowledge, it was the first time that an multimodal LLM-powered self-evolving VLN framework was proposed. Specifically, SE-VLN comprised three core modules, i.e., a hierarchical memory module to transfer successful and failure cases into reusable knowledge, a retrieval-augmented thought-based reasoning module to retrieve experience and enable multi-step decision-making, and a reflection module to realize continual evolution. Comprehensive tests illustrated that the SE-VLN achieved navigation success rates of 57% and 35.2% in unseen environments, representing relative performance improvements of 23.9% and 15.0% over current state-of-the-art methods on R2R and REVERSE datasets, respectively. Moreover, the SE-VLN showed performance improvement with increasing experience repository, elucidating its great potential as a self-evolving agent framework for VLN.
@@ Introduction
Vision-language navigation (VLN), as a key technology connecting human natural language and robot visual navigation, aims to enable embodied agents to autonomously plan paths and complete visual navigation tasks in unseen environments based on human language instructions (Anderson et al. 2018). Despite its rapid development (Chen et al. 2024b; Schumann et al. 2024; Li et al. 2024; Lin et al. 2024), especially with the help of emerging LLMs (Zhou, Hong, and Wu 2024; Chen et al. 2024a; Li, Chen, and Lin 2024; Long et al. 2024), current VLN methods lack the advanced navigation capabilities of natural agents (e.g., bats, migratory birds, humpback whales) and shows critical shortcomings in autonomous evolution through experience. In fact, an experienced horse refines its neural circuitry through repeated journeys. Its navigation skill distills wisdom from experience and evolves autonomously through adaptation, providing a biological blueprint for existing VLN to overcome bottlenecks of ”data dependency” and ”scene generalization”.
Current VLN methods (Li, Chen, and Lin 2024; Zhou, Hong, and Wu 2024; Chen et al. 2024a) had significant limitations in terms of static experience utilization. They treated historical trajectory data as a fixed replay buffer, only using it to maintain decision consistency for the current task, and failed to extract reusable general knowledge from dynamic processes. Another drawback of existing VLN methods (Chen et al. 2024a; Zhou, Hong, and Wu 2024) lay in their limited reasoning capabilities. They relied on the static knowledge invocation of pre-trained models and failed to effectively integrate experience for multi-step decision-making. This stood in stark contrast to the biological intelligence of natural agents such as horses, which adjusted navigation paths by combining environmental memories with real-time perceptions, helping gain precise alignment between visual perceptions and language instructions when processing complex commands. Furthermore, the ”evolution” processes of existing VLN agents (Anderson et al. 2018; Lin et al. 2024, 2025) ware highly dependent on manual hyperparameter tuning and model iteration, which was completely different from the mechanism which biological agents used to achieve self-evolution through natural selection and neural plasticity.
@@ A Reflection Module
As shown in Fig.4, we examined the evolution ability of the VLN framework by adjusting the number of experience entries (set to 0, 10, 30, and 50). As the number of experience increased, the overall performance of SE-VLN showed an improvement trend, i.e., OSR improved from 64.1% to 68.0%, and both SR and SPL were also improved with experience. These indicated that SE-VLN could effectively utilize past experience for continual evolution across different tasks, thereby progressively improving its navigation performance. Moreover, we observed that as the number of tasks increased, the trend of self-evolution gradually flattened out. This might have been due to the limitations of LLMs’ reasoning capabilities, and the navigation experience would gradually tend toward homogenization.
@@
Existing methods were categorized into two types: supervised learning approaches and training-free approaches. As shown in Table 2, we evaluated the navigation performance across various scenarios using a sampled subset of 72 scenes from the R2R dataset. To ensure a fair comparison, we utilized GPT-4o as the inference model for testing both methods. Our results demonstrated that SE-VLN, significantly outperformed MapGPT, the SOTA LLM-powered VLN method. Specifically, the SE-VLN surpassed MapGPT (GPT-4o based) by 0.55 meters in NE, 7.2% in OSR, 8.4% in SR, and 9.2% in SPL. This validated the effectiveness of our approach in improving the navigation performance of LLM-powered agents without the need for training on large-scale datasets.
Furthermore, we conducted additional comparisons of the navigation performance between SE-VLN and previous models on a larger validation set that included 11 scenes with 783 trajectories. As shown in Table 3, despite the distributional differences, SE-VLN exhibited slight improvements across key metrics compared to the results from the 72-scene experiments, with further improvements of SR and SPL by 2.3% and 3.0%, respectively. This might be attributed to the fewer number of scenes in the validation set, where the experience repository might contain similar perceptual and decision experience. This result indicated that the experience repository mechanism of the SE-VLN could effectively capture the decision-making logic across tasks. When there were task patterns in new tasks that were similar to those in the experience repository, the agent could quickly transfer and reuse relevant experience, thereby offsetting the performance loss caused by distribution differences.
@@ Conclusion
Figure 5: Performance of Different MLLMs on the R2R Sampling Subsets.
@@
Figure 9: The prompts for the retrieval-augmented thought-based reasoning module consisted of two parts: the prefix prompt and the input prompt. These respectively provided the prerequisite information needed for the task and specified the input required at each step.